OpenStudy (anonymous):

Will give medal! Please help! Mrs. Miller sells a house for $179,000. If she earns a commission of 6%, how much money does she earn? Write a proportion and show your work. Can someone just help me with writing the proportion? Thanks!! :)

OpenStudy (anonymous):

179,000/100 = x/6 Cross multiply. 179,000 * 6 = 100 * x 1074000 = 100x 10,740 = x <--- Answer. (:

OpenStudy (anonymous):

Thanks so much! You didn't have to solve it, but thanks for trying :)

OpenStudy (anonymous):

You're welcome. (: